### v1

# Can a $\xi R\phi^{2}$ non-minimal coupling contribute to the non-relativistic gravitational potential?

Abstract
We investigate the contribution of the $\sqrt{-g}\xi R\phi^{2}$ non-minimal interaction for a massive scalar to the gravitational potential, found via the non-relativistic 2-2 scattering with one and two graviton exchanges up to one loop order. Such coupling can be most naturally motivated from the renormalisation of a scalar field theory with quartic self interaction in a curved spacetime. This coupling is qualitatively different from the minimal ones, like $\sim\sqrt{G}h^{\mu\nu}T_{\mu\nu}$, as the vertices corresponding to the former does not explicitly contain any scalar momenta, but instead explicitly contains the momentum of the graviton. We show that the insertion of one/two graviton-two scalar non-minimal vertices in the standard 2-2 scattering Feynman diagrams does not yield any contribution for the potential up to one loop, owing to the aforesaid explicit appearance of graviton’s transfer momentum in the numerator of the Feynman amplitude. In order to complement this effect, the most natural choice seems to consider the three graviton vertex generated by the $\sim\Lambda\sqrt{-g}/G$ term in the action, where $\Lambda$ is the cosmological constant. This vertex does not contain explicitly any graviton momentum. With this, and assuming short scale scattering much small compared to the Hubble horizon, we compute the seagull, the vacuum polarisation and the fish diagrams and obtain the non-relativistic scattering amplitudes. They indeed yield a two body gravitational potential at ${\cal O}(\xi\Lambda G^{2})$. We further point out some qualitative differences of this potential with that of the standard result of $\xi=0$.
Keywords : $\xi R\phi^{2}$ interaction, perturbative quantum gravity, cosmological constant, gravitational potential
1 Introduction
There has been a lot of effort in the past and present to understand the low energy effects of perturbative quantum gravity, even though it is not renormalisable. The motivation behind this is to see its predictions at first few order of the perturbation theory, with the hope that they will be verifiable in a not too far away future. We also hope that such perturbative framework will be smoothly embedded in a complete theory of quantum gravity one day. In particular, the perturbative quantum gravity might tell us whether at all the concept of graviton as a quantum massless spin-2 field is meaningful. We refer our reader to [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and references therein for various tree and one loop computations in this framework, chiefly focussing on the leading quantum correction to the gravitational or Newton potential, emerging from the graviton exchange of two non-relativistic particles. See also [17, 18, 19] and references therein for computation of bending of light. We also refer our reader to [20, 21, 22, 23] and references therein for computation of the corrections to Newton potential via effective field theory techniques. Moreover, see [24, 25, 26, 27, 28] for inclusion of a cosmological constant at small scales, where the background metric can still be taken to be $\eta_{\mu\nu}$. Finally, see e.g. [29, 30, 31, 32, 33] and references therein for discussions in curved spacetimes like de Sitter.
In this work we wish to investigate the curvature-scalar non-minimal coupling term, $\xi\sqrt{-g}R\phi^{2}$, 2, in the context of perturbative quantum gravity for a massive scalar field, with $\xi$ being the non-minimal coupling parameter. Such coupling naturally arises in the presence of background curvature while renormalising a scalar field theory with quartic self interaction [34]. What will be the contribution of this term at lowest orders in the non-minimal coupling $\xi$ and at one loop, to the correction of the Newton potential between two massive scalar particles? Note that this coupling is qualitatively different from the minimal ones. This is because for a scalar field theory with quadratic kinetic term, in the matter-graviton interaction there can be no derivative of the graviton (such as $\kappa h^{\mu\nu}T_{\mu\nu}$), whereas for the non-minimal interaction, there is no derivative of the scalar at all (6). Due to the latter, we will see that replacing any standard matter-graviton vertex by an ${\cal O}(\xi)$ one in any 2-2 scattering diagram, yields no contribution for the potential at tree or one loop level, owing to the appearance of graviton’s transfer momentum in the numerator of the Feynman amplitude. As we need to consider the two scalar-one graviton and two scalar-two graviton vertices generated by the $\sqrt{-g}\xi R\phi^{2}$ term, the diagrams will have the same topology as that of [3, 4]. Thus in order to find out whether at all this interaction can make any contribution to the non-relativistic potential, we must check whether there is any possibility to ‘nullify’ the effect of the aforementioned derivative or momentum of the graviton. We will see that this is possible in the Einstein gravity, if we consider a cosmological constant and the corresponding ${\cal O}(\Lambda\kappa)$ three-graviton vertex generated by the $4\Lambda\sqrt{-g}/\kappa^{2}$ term, 2. Unlike the usual three graviton vertex generated by $2\sqrt{-g}R/\kappa^{2}$, there will be no derivative of the graviton here. Accordingly, we will see in 3 that the seagull (2), the vacuum polarisation (3) and the fish diagrams (4) will contribute to the gravitational potential, at ${\cal O}(\xi\Lambda\kappa^{4})$. We will further emphasise qualitative differences of the potential thus found from the standard ones ($\xi=0=\Lambda$). We also recall that for a conformally symmetric scalar field ($m^{2}=0,\ \xi=1/6$), static spherically symmetric black hole solution is known to exist since long [35, 36, 37, 38]. To the best of our knowledge, the $\xi\sqrt{-g}R\phi^{2}$ non-minimal coupling is rather less investigated from the quantum gravity perspective. Due to its aforementioned qualitatively different nature from the standard matter-graviton coupling, it seems to be an important job to probe the effect of the first on quantum gravity predictions.
We also note that in general the presence of a $\Lambda$ prevents us from doing quantum field theory in a flat background, and instead one should do it in a de Sitter background, which reads for example in the static coordinatisation,
$ds^{2}=-fdt^{2}+f^{-1}dr^{2}+r^{2}d\Omega^{2},$ (1)
where $f=1-\Lambda r^{2}/3$. However much inside the Hubble horizon ($r\ll\sqrt{3/\Lambda}$), the spacetime can be split into a Minkowski background plus a perturbation. The Schwarzschild-de Sitter spacetime, describing the gravitational filed of a static mass inside the de Sitter universe, is given by $f(r)=1-2MG/r-\Lambda r^{2}/3$ in 1. The region with smaller $r$ values are dominated by the attraction due to the mass $M$, and the larger ones by the repulsion due to $\Lambda$. The attraction decreases monotonically with increasing $r$ whereas the repulsion increases. Accordingly, there is a radius, $R_{\rm TA,max}=(3MG/\Lambda)^{1/3},$ known as the maximum turn around radius, around which the attraction and repulsion balance each other, e.g. [39]. It is clear that we may do a decomposition of flat spacetime plus perturbation inside this radius, to look into short wavelength phenomena. Such decomposition in the presence of a $\Lambda$ has been done earlier in [24, 25, 26, 27, 28].
Before we go into the computations, we will summarise below the basic technical ingredients we will need. We will work with mostly positive signature of the metric in four spacetime dimensions. Our notation for symmetrisation will be : $X_{(\mu\nu)}=X_{\mu\nu}+X_{\nu\mu}$.
2 The basic ingredients
The action reads
$S=\frac{2}{\kappa^{2}}\int d^{4}x\sqrt{-g}\ \left(R-2\Lambda\right)-\frac{1}{2}\int d^{4}x\sqrt{-g}\ \left[g^{\mu\nu}(\partial_{\mu}\phi)(\partial_{\nu}\phi)+m^{2}\phi^{2}+\xi R\phi^{2}\right]$ (2)
where $\kappa^{2}=32\pi G$. We wish to treat the non-minimal interaction term perturbatively. The metric is broken into background plus perturbation
$g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu};\qquad g^{\mu\nu}=\eta^{\mu\nu}-\kappa h^{\mu\nu}+\kappa^{2}h^{\mu}{}_{\alpha}h^{\alpha\nu}+\cdots$ (3)
Using the De Donder gauge,
$\partial_{\mu}\left(h^{\mu}{}_{\nu}-\frac{1}{2}\delta^{\mu}_{\nu}h\right)=0,$
the leading, free and linearised Einstein equation reads
$\partial^{2}\left(h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}\right)-\frac{2\Lambda}{\kappa}\eta_{\mu\nu}=0,$ (4)
giving
$h_{\mu\nu}=h_{\mu\nu}|_{\rm hom}-\frac{\Lambda r^{2}}{3\kappa}\eta_{\mu\nu},$ (5)
where the first term on the right hand side is the usual homogeneous massless spin-2 plane wave solution which we will use for field quantisation, whereas the second term corresponds to the linearised, classical de Sitter perturbation yielding the background curvature, $R=4\Lambda$. The field quantisation with respect to $h_{\mu\nu}|_{\rm hom}$ yields the graviton propagator. We will restrict ourselves to ${\cal O}(\Lambda)$ only. We now make the following leading expansions for the scalar and the $\Lambda$ part of the action,
$S_{\phi}=-\frac{1}{2}\int d^{4}x\left[(\partial\phi)^{2}+m^{2}\phi^{2}\right]+\frac{\kappa}{2}\int d^{4}xh^{\mu\nu}\left[(\partial_{\mu}\phi)(\partial_{\nu}\phi)-\frac{1}{2}\eta_{\mu\nu}\left((\partial\phi)^{2}+m^{2}\phi^{2}\right)\right]$
$-\frac{\kappa^{2}}{2}\int d^{4}x\left[\left(h^{\mu}{}_{\alpha}h^{\alpha\nu}-\frac{1}{2}hh^{\mu\nu}+\frac{h^{2}}{8}\eta^{\mu\nu}-\frac{1}{4}h_{\alpha\beta}h^{\alpha\beta}\eta^{\mu\nu}\right)(\partial_{\mu}\phi)(\partial_{\nu}\phi)+\frac{m^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)\phi^{2}\right]$
$+\frac{\xi}{2}\int d^{4}x\left[\kappa\phi^{2}\partial^{2}h-\kappa^{2}\phi^{2}\left\{-\frac{1}{4}h\partial^{2}h-\frac{1}{2}(\partial^{\lambda}h^{\mu\alpha})(\partial_{\alpha}h_{\mu\lambda})+h^{\mu\nu}\partial^{2}h_{\mu\nu}+\frac{3}{4}(\partial^{\lambda}h^{\mu\nu})(\partial_{\lambda}h_{\mu\nu})\right\}\right]+\ {\cal O}(\kappa^{3},\xi\kappa^{3}),$
$S_{\Lambda}=-\frac{4\Lambda}{\kappa^{2}}\int d^{d}x\left[1+\frac{\kappa h}{2}+\frac{\kappa^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)+\frac{\kappa^{3}}{2}\left(\frac{1}{3}h_{\mu\nu}h^{\nu\lambda}h_{\lambda}{}^{\mu}-\frac{1}{4}hh_{\mu\nu}h^{\mu\nu}+\frac{h^{3}}{24}\right)\right]+\ {\cal O}(\Lambda\kappa^{2}),$ (6)
where we have used $m^{2}\gg\Lambda$, i.e. the Compton wavelength of the scalar is much small compared to the Hubble horizon, as required. The scalar and the graviton propagators read
$\Delta(k)=-\frac{1}{k^{2}+m^{2}};\qquad\Delta_{\mu\nu\alpha\beta}(k)=-\frac{P_{\mu\nu\alpha\beta}}{k^{2}},\qquad P_{\mu\nu\alpha\beta}=\frac{1}{2}\left(\eta_{\mu\alpha}\eta_{\nu\beta}+\eta_{\mu\beta}\eta_{\nu\alpha}-\eta_{\mu\nu}\eta_{\alpha\beta}\right),$ (7)
where $\mu,\nu$ and $\alpha,\beta$ pairwise correspond to two graviton fields, $h_{\mu\nu}$ and $h_{\alpha\beta}$. Note in 6 that there is a term quadratic in $h_{\mu\nu}$ in $S_{\Lambda}$, which could also make a contribution proportional to $\Lambda h_{\mu\nu}$ in 4. However, this is subleading compared to the $\eta_{\mu\nu}$ term. Note also from 5 that the classical term will make contribution beyond linear order in $\Lambda$ in $S_{\Lambda}$, and hence we will take in our calculations $h_{\mu\nu}$ to be purely quantum in the cubic interaction. This term will generate a three graviton vertex of ${\cal O}(\kappa\Lambda)$ with no explicit appearance of the graviton momentum. The usual three graviton vertex originating from the $\sqrt{-g}R/2\kappa^{2}$ term (with momenta $l$, $q$, lql-q along the three lines) reads, e.g. [4],
$V^{(3)\mu\nu}_{\alpha\beta\gamma\delta}(\kappa;l,q)$ $=-\frac{i\kappa}{2}\left[P_{\alpha\beta\gamma\delta}\left(l^{\mu}l^{\nu}+(l-q)^{\mu}(l-q)^{\nu}+q^{\mu}q^{\nu}-\frac{3}{2}\eta^{\mu\nu}q^{2}\right)\right$. (8)
[…]
$\left.+\left\{\left(l^{2}+(l-q)^{2}\right)\left(I_{\alpha\beta}{}^{\mu\sigma}I_{\gamma\delta\sigma}{}^{\nu}+I_{\gamma\delta}{}^{\mu\sigma}I_{\alpha\beta\sigma}{}^{\nu}-\frac{1}{2}\eta^{\mu\nu}{\cal P}_{\alpha\beta\gamma\delta}\right)-\left(I_{\gamma\delta}{}^{\mu\nu}\eta_{\alpha\beta}l^{2}+I_{\alpha\beta}{}^{\mu\nu}\eta_{\gamma\delta}(l-q)^{2}\right)\right\}\right],\qquad$
with
$I_{\mu\nu\lambda\rho}=\frac{1}{2}\left(\eta_{\mu\lambda}\eta_{\nu\rho}+\eta_{\mu\rho}\eta_{\nu\lambda}\right).$
We will also require the standard integral [40]
$I^{r}_{s}=\int\frac{d^{d}k}{(2\pi)^{d}}\frac{(k^{2})^{r}}{(k^{2}-a^{2})^{s}}=\frac{i(-1)^{r+s}}{(4\pi)^{d/2}}\frac{1}{(a^{2})^{s-r-d/2}}\frac{\Gamma(r+d/2)\Gamma(s-r-d/2)}{\Gamma(d/2)\Gamma(s)}.$ (9)
Let us first see how the non-minimal term contributes to the gravitational interaction, without any $\Lambda$. The simplest process is given by the first of 1. From 6, we see that the non-minimal one graviton - two scalar vertex function reads : $i\xi\kappa(-q^{2})\eta_{\mu\nu}/2$. The $k_{2}-k^{\prime}_{2}$ vertex is the usual minimal one graviton - two scalar vertex. The Feynman amplitude reads,
$i{\cal M}=-i.i\kappa\frac{i\xi\kappa(-q^{2})\eta_{\mu\nu}P^{\mu\nu\alpha\beta}\left(k_{1(\alpha}k^{\prime}_{1\beta)}-\eta_{\alpha\beta}(k_{1}\cdot k^{\prime}_{1}+m_{1}^{2})\right)}{2^{2}q^{2}}$ (10)
[…]
Figure 1: The tree and the box diagrams for 2-2 scattering with graviton exchanges. Solid and wavy lines respectively denote scalar and graviton propagators. The thick circle denotes non-minimal vertex of ${\cal O}(\xi\kappa)$. The external momenta are on shell, ${k^{2}_{1}}={k^{\prime}}_{1}^{2}=-m_{1}^{2}$, ${k^{2}_{2}}={k^{\prime}}_{2}^{2}=-m_{2}^{2}$. None of the diagrams yields any non-analytic contribution in the transfer momentum, $q=k_{1}-k^{\prime}_{1}=k^{\prime}_{2}-k_{2}$, and hence to gravitational potential. Similar conclusion holds for the crossed box and other one loop diagrams in the absence of any $\Lambda$-three graviton vertex. See main text for detail.
3 The $\Lambda$-three graviton vertex and the gravitational potential at ${\cal O}(\xi\kappa^{4}\Lambda)$
Figure 2: The seagull diagram for 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$, with the $\Lambda$-three graviton vertex. There are three more diagrams, found by the interchange, $(k_{1},k^{\prime}_{1})\ \leftrightarrow\ (k_{2},k^{\prime}_{2})$.
Figure 3: The vacuum polarisation diagram for the 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$. One of the cubic graviton vertex corresponds to the $\Lambda$-three graviton vertex, 6, whereas the other one is the usual, 8. Also as earlier, the thick circle denotes non-minimal vertex.
Figure 4: The fish diagram for the 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$. The cubic vertex is the $\Lambda$-three graviton vertex.
3.1 The seagull
There are total six diagrams for the seagull class, 2. Using 6 for $S_{\Lambda}$, the first of them gives the contribution ($d=4-\epsilon$)
$-i{\cal M}^{\rm 2.1}=\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{3}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda\rho}\eta_{\gamma\delta}P^{\lambda\rho\lambda^{\prime}\rho^{\prime}}P^{\gamma\delta\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\lambda^{\prime}}(l-k^{\prime}_{2})_{\rho^{\prime})}-\eta_{\lambda^{\prime}\rho^{\prime}}(k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}$
$\times\frac{\left\{k_{2(\gamma^{\prime}}(l-k_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$-\frac{\Lambda\kappa^{4}\xi}{2^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda^{\prime}\mu^{\prime}}\eta_{\nu^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\mu^{\prime}\nu^{\prime}\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\gamma^{\prime}}(l-k^{\prime}_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}$
$\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
[…]
$\times\frac{\left\{k_{2(\gamma}(l-k_{2})_{\delta)}-\eta_{\gamma\delta}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+\Lambda\kappa^{4}\xi\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\alpha\beta}P_{\alpha\beta\mu\nu}P^{\nu\lambda\gamma\delta}P_{\lambda}{}^{\mu\alpha^{\prime}\beta^{\prime}}\left\{k^{\prime}_{2(\gamma}(l-k^{\prime}_{2})_{\delta)}-\eta_{\gamma\delta}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$=-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k^{\prime}_{2}\cdot(l-k^{\prime}_{2})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}=0,$ (11)
where $\mu$ is the mass scale for renormalisation. Likewise if we interchange pairwise $k_{1},k^{\prime}_{1}$ with $k_{2},k^{\prime}_{2}$, we obtain vanishing contribution.
We next place the non-minimal vertex on the $k_{2}^{\prime}$ momentum, the second of 2. Using the on-shell and non-relativistic relations ($k_{1}^{2}={k^{\prime}_{1}}^{2}=-m_{1}^{2}$, $k_{2}^{2}={k^{\prime}_{2}}^{2}=-m_{2}^{2}$, $k_{1}\cdot k_{2}\approx-m_{1}m_{2},\,k^{\prime}_{1}\cdot k^{\prime}_{2}\approx-m_{1}m_{2}$), we have
$-i{\cal M}^{\rm 2.2}=\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{4}q^{2}}k_{1(\mu}k^{\prime}_{1\nu)}\eta^{\alpha\beta}\eta^{\gamma\delta}\eta^{\lambda\rho}\eta^{\lambda^{\prime}\rho^{\prime}}P^{\mu\nu}{}_{\alpha\beta}P_{\lambda\rho\lambda^{\prime}\rho^{\prime}}P_{\gamma\delta}{}^{\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\alpha^{\prime}}(l+k_{2})_{\beta^{\prime})}-l\cdot k_{2}\ \eta_{\alpha^{\prime}\beta^{\prime}}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}k_{1(\mu}k^{\prime}_{1\nu)}P^{\mu\nu}{}_{\mu^{\prime}\nu^{\prime}}\eta^{\gamma\delta}P^{\nu^{\prime}\lambda^{\prime}}{}_{\gamma\delta}P_{\lambda^{\prime}}{}^{\mu^{\prime}\alpha\beta}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\alpha}(l+k_{2})_{\beta)}-l\cdot k_{2}\ \eta_{\alpha\beta}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{3}q^{2}}k_{1(\alpha}k^{\prime}_{1\beta)}\eta^{\lambda^{\prime}\rho^{\prime}}P^{\alpha\beta}{}_{\lambda^{\prime}\rho^{\prime}}\eta^{\lambda\rho}P_{\mu\nu\lambda\rho}P^{\mu\nu\gamma\delta}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\gamma}(l+k_{2})_{\delta)}-l\cdot k_{2}\ \eta_{\gamma\delta}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{2}q^{2}}k_{1(\alpha}k^{\prime}_{1\beta)}P^{\alpha\beta}{}_{\mu\nu}\eta^{\lambda\rho}P_{\lambda\rho\lambda^{\prime}\rho^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}P^{\mu\nu\gamma\delta}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\gamma}(l+k_{2})_{\delta)}-l\cdot k_{2}\ \eta_{\gamma\delta}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$=-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$=-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})},$ (12)
which we evaluate using 9. The divergence can be absorbed in the renormalisation counterterm for the one graviton-two scalar vertex diagram embedded in 2. The part relevant to the potential (i.e., non-analytic in the transfer momentum $q^{2}$) reads
$-i{\cal M}^{\rm 2.2}\equiv-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{8\pi^{2}q^{2}}$ (13)
If we place the non-minimal vertex on the $k_{2}$ vertex, we obtain exactly the same result as above. Likewise, if we pairwise interchange the $k_{1},k_{1}^{\prime}$ and $k_{2},k_{2}^{\prime}$ vertices, we obtain the same result, Thus we have the total contribution to the gravitational potential for the seagull class of diagrams, 2,
$-i{\cal M}^{\rm seagull}_{\rm tot.}\equiv-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}q^{2}}$ (14)
3.2 The vacuum polarisation
The vacuum polarisation class of diagrams are given by 3. There are two three graviton vertices involved here – one is associated with $\Lambda$, whereas the other one is the usual, 8. Using $k_{1}\cdot q\approx q^{2}/2$, $k_{1}^{\prime}\cdot q\approx-q^{2}/2$, $k_{2}\cdot q\approx-q^{2}/2$ and $k^{\prime}_{2}\cdot q\approx q^{2}/2$, we compute for the first diagram
$-i{\cal M}^{\rm 3.1}=\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{8}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}\eta^{\gamma^{\prime}\delta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P_{\gamma^{\prime}\delta^{\prime}\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$+3\mu^{4\epsilon}\kappa^{4}\Lambda\xi(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\beta^{\prime}\gamma^{\prime}}{}_{\lambda\rho}P_{\gamma^{\prime}}{}^{\alpha^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{4}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{2}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\rho}\eta_{\lambda^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}=0,$ (15)
where $\bar{\tau}^{\mu\nu\lambda\rho\gamma\delta}$ is the three graviton vertex of 8, modulo the factor $i\kappa/2$. Likewise, if the $k_{2}-k_{2}^{\prime}$ is non-minimal and it is connected to the $\Lambda$-three graviton vertex (i.e., the third of 3), the contribution vanishes. Similarly, we find all the contributions for this class of diagrams vanish.
3.3 The fish diagram
This class of diagrams is given by 4, and the cubic vertex here is the $\Lambda$-three graviton vertex. Let us first place the non-minimal vertex in the $k_{1}-k_{1}^{\prime}$ vertex. The two scalar-two graviton vertex is the usual (${\cal O}(\kappa)$). We thus have for the first diagram,
$-i{\cal M}^{\rm 4.1}=-\Lambda\mu^{4\epsilon}\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\lambda^{\prime}\rho^{\prime}}}{q^{2}}\eta^{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\eta_{\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\eta}\eta_{\rho^{\prime}\sigma^{\prime}}P^{\rho^{\prime}\sigma^{\prime}}{}_{\rho\sigma}}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$-8\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\nu^{\prime}\lambda^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}}{}^{\mu^{\prime}}_{\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$+2\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}\eta^{\mu^{\prime}\nu^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$+4\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda\eta}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$ (16)
where we have abbreviated for the two scalar-two graviton vertex [4],
$\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}={\tau}_{2,2}^{\lambda\eta\rho\sigma}/\kappa^{2}=\left[\left\{I^{\eta\lambda\alpha\delta}I^{\rho\sigma\beta}{}_{\delta}-\frac{1}{4}\left(\eta^{\eta\lambda}I^{\rho\sigma\alpha\beta}+\eta^{\rho\sigma}I^{\eta\lambda\alpha\beta}\right)\right\}k_{2(\alpha}k^{\prime}_{2\beta)}-\frac{1}{2}\left(I^{\eta\lambda\rho\sigma}-\frac{1}{2}\eta^{\eta\lambda}\eta^{\rho\sigma}\right)(k_{2}\cdot k^{\prime}_{2}+m_{2}^{2})\right]$ (17)
A little algebra shows that 16 is vanishing. Likewise if the ${\cal O}(\kappa)$ non-minimal vertex is $k_{2}-k^{\prime}_{2}$, we find vanishing contribution (the third of 4).
We next come to the case when the 2-scalar 2-graviton vertex is non-minimal, ${\cal O}(\xi\kappa^{2})$. We have for the second of 4,
$-i{\cal M}^{\rm 4.2}=\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{4}k_{1(\mu}k^{\prime}_{1\nu)}P^{\mu\nu\mu^{\prime}\nu^{\prime}}\eta_{\mu^{\prime}\nu^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\alpha}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\nu}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\nu}\right\}$
$+2\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P^{\rho^{\prime}\gamma^{\prime}\mu\alpha}{P_{\gamma^{\prime}}{}^{\lambda^{\prime}}}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P^{\gamma^{\prime}\rho^{\prime}\mu\nu}P^{\lambda^{\prime}}{}_{\gamma^{\prime}\mu\nu}\right\}$
$-\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\alpha}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\nu}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\nu}\right\}$
$-\mu^{3\epsilon}\frac{\Lambda\kappa^{4}\xi}{2}k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta_{\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P_{\gamma\delta}{}^{\mu\alpha}P^{\gamma\delta}{}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\gamma\delta}{}^{\mu\nu}P^{\gamma\delta}{}_{\mu\nu}\right\}$
$=\frac{\mu^{3\epsilon}m_{1}^{2}\Lambda\kappa^{4}\xi}{4}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}-\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2k_{1\mu}k^{\prime}_{1\nu}l^{\mu}l^{\nu}+q^{2}l\cdot q/2}{l^{2}(l-q)^{2}}$ (18)
Only the first integral contributes to the potential at leading order, which is $\sim\ln q^{2}$. Likewise by replacing $m_{1}$ by $m_{2}$, we obtain the result if $k_{2}-k^{\prime}_{2}$ is the non-minimal two scalar-two graviton vertex. Putting things together now including 14, we have the total Feynman amplitude for 2-2 scattering at ${\cal O}(\xi\Lambda\kappa^{4})$, in the non-relativistic limit,
$-i{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}=-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}\vec{q}^{2}}-\frac{i(m_{1}^{2}+m_{2}^{2})\Lambda\kappa^{4}\xi}{4\pi^{2}}\ln\frac{\vec{q}^{2}}{\mu^{2}}$ (19)
4 The final result and discussion
The non-relativistic gravitational potential in the first Born approximation is defined as [3, 4],
$V(\vec{r})=-\frac{1}{4m_{1}m_{2}}\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot\vec{r}}{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}(\vec{q}^{2})$
Using the above, we obtain from 19 the leading contribution of the non-minimal coupling parameter $\xi$ to the two body gravitational potential,
$V(G^{2},\Lambda,\xi,r)=-\frac{32G^{2}\Lambda\xi m_{1}m_{2}}{\pi r}\left[1-\frac{m_{1}^{2}+m_{2}^{2}}{m^{2}_{1}m^{2}_{2}}\frac{\mu l\hbar}{cr^{2}}\right],$ (20)
where the factor $\mu l\hbar/c$ ensures dimensional consistency, with $l$ being a length scale. In the absence of a cosmological constant, there can be no contribution from $\xi$ to the potential, at least to this order of the perturbation theory. Note that while the right hand side of the momentum space expression 19 is perfectly dimensionally consistent, the Fourier transform of the logarithm necessitates the insertion of such factor. How do we fix $l$? For example, we may set $l=R_{\rm TA,max}$, i.e. the maximum turn around radius defined below 1, as one of the natural length scales in this context, beyond which this result cannot be trusted. However, this choice is not unique and anything concrete beyond this remains elusive to us so far. Note that such ambiguity is absent for the $\xi=0$ part of the potential, due to the appearance of different powers of $G$ in different terms, fixing such term uniquely [3, 4]. (Here we do not have any term first order in $G$).
Note that the second term in the potential diverges if either of the masses is vanishing. If we understand it correctly, we do not expect such behaviour from a non-relativistic quantity. This problem originates from the absence of the scalar rest mass term in the scalar-graviton non-minimal vertex, 6. Does this imply that the $\ln\vec{q}^{2}$ term of 19 should not be related to a physical gravitational potential, or, we simply accept 20 as it is? Once again, any clear reasoning in the favour of either of this remains elusive to us so far.
Further note that the potential of 20, even though corresponds to pure short scale physics, contains two fundamental length scales (apart from $R_{\rm TA,max}$) — the shortest or Planck length corresponding to $G$, and the longest or Hubble horizon scale corresponding to $\Lambda$. We believe this is a novel feature encoded in a short scale physics computation.
It seems an important task to further investigate the effect of $\xi$, as well as the vertex associated with $\Lambda$ on various perturbative quantum gravity phenomenon. As we have emphasised, this coupling is qualitatively different from the minimal ones. It would thus be interesting to investigate the effect of $\xi$ on self energy, various matter-graviton vertices and RG perspectives. A look into the soft theorems might also be interesting in this context. We wish to come to these issues in our future publications.
Acknowledgements : SB would like to thank Amitabha Lahiri and Parthasarathi Majumdar for useful discussions.